# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

### v1

@@ Abstract
Pre-trained machine learning (ML) predictions have been increasingly used to complement incomplete data to enable downstream scientific inquiries, but their naive integration risks biased inferences. Recently, multiple methods have been developed to provide valid inference with ML imputations regardless of prediction quality and to enhance efficiency relative to complete-case analyses. However, existing approaches are often limited to missing outcomes under a missing-completely-at-random (MCAR) assumption, failing to handle general missingness patterns under the more realistic missing-at-random (MAR) assumption. This paper develops a novel method which delivers valid statistical inference framework for general Z-estimation problems using ML imputations under the MAR assumption and for general missingness patterns. The core technical idea is to stratify observations by distinct missingness patterns and construct an estimator by appropriately weighting and aggregating pattern-specific information through a masking-and-imputation procedure on the complete cases. We provide theoretical guarantees of asymptotic normality of the proposed estimator and efficiency dominance over weighted complete-case analyses. Practically, the method affords simple implementations by leveraging existing weighted complete-case analysis software. Extensive simulations are carried out to validate theoretical results. The paper concludes with a brief discussion on practical implications, limitations, and potential future directions.
@@ 1 Introduction
With the rapid development of machine learning (ML) and artificial intelligence (AI) techniques, ML models, particularly those trained on large-scale datasets, have become increasingly capable of making highly accurate predictions. These predictions can be treated as imputations for missing data (in both dependent $Y$ and independent variables $\mathbf{X}$) due to nonresponse, expensive data collection costs, or other factors. Because ML-imputed datasets can be seamlessly applied in downstream statistical analyses, such appealing impute-then-analyze procedures is a double-edged sword. On the one hand, the ease of obtaining ML predictions for imputation catalyzes wide adoptions across various disciplines, including finance (Bryzgalova et al.,, 2025), social science (Egami et al.,, 2023), clinical psychology (Little,, 2024), biomedicine (Barrabés et al.,, 2024), and public health (Perkins et al.,, 2018) studies, where missing data appear commonly during the data collection process. On the other hand, imputing missing values with ML predictions of arbitrary quality may produce datasets that have different conditional and marginal distributions from the underlying true distribution. Proceeding with analysis without careful consideration may result in biased estimation and invalid inference.
@@
To address this issue, recent research has explored prediction-based methods to incorporate ML predictions for single imputation while guaranteeing valid statistical inference on datasets with partially missing outcomes (see, for example, Wang et al., (2020); Angelopoulos et al., 2023a ; Egami et al., (2023); Salerno et al., 2025a , or Salerno et al., 2025b ). Specifically, prediction-powered inference (PPI, Angelopoulos et al., 2023a, ) is a framework where inference is conducted using ML-imputed outcomes, with model parameter estimates calibrated, or “rectified”, using a complete-case subset. Since this approach does not impose assumptions on the imputation model itself, it allows for a broad selection of ML-based imputation methods, the quality of which impacts the efficiency but not the validity of inference (e.g., consistency, achieving nominal coverage rate). Several extensions of this framework have adapted the idea of PPI to important areas including causal inference (Demirel et al.,, 2024), Bayesian inference (Hofer et al.,, 2024; Ročková and O’Hagan,, 2025), and genome-wide association studies (GWAS, Miao et al.,, 2024). Furthermore, recognizing that the PPI approach does not guarantee a smaller variance relative to complete-case analysis (CCA), a few authors have proposed weighting the rectifier with tuning parameters to achieve better efficiency than CCA (Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan et al.,, 2024; Gronsbell et al.,, 2024; Ji et al.,, 2025).
@@
Despite the tremendous progress and development in prediction-based inference, two key challenges remain unaddressed. First, the existing literature considers only the assumption of missing completely at random (MCAR), leaving open the question of the applicability of prediction-based inference under the more common and complicated assumption, that is, the assumption missing at random (MAR) (Rubin,, 1976; Little and Rubin,, 2019). Except for several specific cases such as linear regression analysis and odds ratio estimation under certain missing probability assumptions  (Example 3.3 and 3.4, Little and Rubin,, 2019), failure to account for the MAR assumption can result in biased estimates. Second, the current literature focuses exclusively on partially missing outcomes or specific patterns in covariates missingness (e.g., blockwise, Kluger et al.,, 2025), but does not account for general patterns of non-monotone missingness, where missing values can occur in any cell within the tabular (“rectangular”) dataset. For example, in regression analysis, missing values are common in both outcomes and covariates, and we may observe different patterns of missingness in different subjects, e.g., missing $Y$ and $X_{1}$ for one subject and missing $Y$ and $X_{2}$ for the other.
@@
In this paper, we address these challenges by proposing a unified framework called Pattern-Stratified PPI (PS-PPI) [^1] By Corollary 3.2, the proposed procedure enjoys efficiency dominance over WCCA (under MAR, general missing patterns). This resembles the property of PPI++ being no less efficient than CCA (under MCAR, outcome missing only). We chose to term our method “PS-PPI” rather than “PS-PPI++” to avoid redundancy. to account for both general missingness patterns and missing-at-random mechanism under the Z-estimation framework. We make a few theoretical and methodological contributions with useful practical implications. First, we generalize the existing PPI-type estimators that only deal with missing outcomes under MCAR. We develop an inferential procedure that integrates the concept of missingness patterns that are widely used in semiparametetric inference to deal with non-monotone missing data (Tsiatis,, 2006). We will show the proposed estimator belongs to the class of Augmented Inverse Probability Weighted (AIPW) estimators, building on a prior work under MCAR with only outcomes subject to missingness (Gronsbell et al.,, 2024). Second, we provide theoretical guarantee that the resulting PS-PPI estimator is at least as efficient as the weighted complete-case analysis (WCCA) even when the imputation models are of low quality. Third, our proposed PS-PPI procedure is straightforward to implement. It works by breaking the analysis into several weighted complete-case (WCC) Z-estimation problems, each of which can be solved with existing statistical packages designed for WCC analysis.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce necessary notation, the concepts of missing data mechanisms and missingness patterns, and the inferential target under the Z-estimation framework; we review the core idea used in existing PPI literature that deals with missing outcome or blockwise missing covariates, along with their limitations. In Section 3, we introduce the details of the proposed PS-PPI procedure and present theoretical results on its asymptotic properties; we provide examples to show how our proposed framework subsumes previous works as special cases. In Section 4, we perform data simulation to verify the validity and superior performance of our proposed PS-PPI relative to alternatives that have limited ability to extract and integrate information from incomplete datasets with multiple different missingness patterns. We discuss the limitations of this work and potential future directions in Section 5.
@@ 2.1 Notation
In this subsection we introduce the notation; a comprehensive glossary of the notation is presented in the Supplementary Materials. Let $\mathcal{D}=\{\bm{Y}_{1},\ldots,\bm{Y}_{N}\}$ denote the full data for $N$ observations, subject to potential missingness, where $\bm{Y}_{i}=(Y_{i1},\ldots,Y_{ip})^{\top}\in{\mathcal{Y}}={\mathcal{Y}}_{1}\times\cdots\times{\mathcal{Y}}_{p}$ represents the $p$-dimensional data for subject $i$, ${\mathcal{Y}}_{j}\subseteq\mathbb{R}$ is the marginal sample space of the $j$-th component in $\bm{Y}_{i}$. Each $\bm{Y}_{i}$ is assumed to be independently and identically sampled from a population distribution $\mathbb{P}_{\bm{Y}}$. In practice, three types of variables may be included in the vector $\bm{Y}_{i}$: (a) a variable treated as the scalar outcome in a specific scientific inquiry; (b) q1q-1 variables treated as covariates; (c) pqp-q auxiliary variables not included in the scientific analysis but potentially useful to predict variables in type (a) and (b). In this paper, unless otherwise noted, we do not explicitly distinguish among these types in our notation to preserve the flexibility of our proposed method in large-scale analyses such as LabWAS (Goldstein et al.,, 2020), where a variable may serve as an outcome in one analysis and as a covariate in another. However, for readers who require this distinction in their scientific inquiries (e.g., via a regression analysis), one may assume that $Y_{i1}$ and $Y_{i2},\ldots,Y_{iq}$} represent the outcome and the covariates, respectively. Note that the partition of $\bm{Y}_{i}$ into three types of variables mentioned above is purely for the ease of understanding and has no impact on the theoretical results in this paper.
We will assume that the analyst may not observe the full data $\mathcal{D}$}. Let $\bm{R}_{i}=(R_{i1},\ldots,R_{ip})^{\top}\\ \in\{0,1\}^{p}$ be a binary vector of $p$ dimension, where $R_{ij}=1$ indicates that $Y_{ij}$ is observed and $R_{ij}=0$ otherwise. Consequently, we denote the observed components $\bm{Y}_{i,\text{obs}}=\{Y_{ij}|R_{ij}=1,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=1\}}{\mathcal{Y}}_{j}$, and similarly the missing components $\bm{Y}_{i,\text{mis}}=\{Y_{ij}|R_{ij}=0,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=0\}}{\mathcal{Y}}_{j}$, for subject $i=1,\ldots,N$. Let ${\mathcal{D}}_{\text{obs}}=\{\bm{Y}_{1,\text{obs}},...,\bm{Y}_{N,\text{obs}}\}$ denote the observed data.
@@ 2.1.1 Missingness patterns and mechanisms
In the existing literature on prediction-based inference, typically only the outcome variable is subject to potential missingness (e.g., Angelopoulos et al., 2023a, ; Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan and Liang,, 2023; Gronsbell et al.,, 2024), with the exception of  Kluger et al., (2025) which focused on blockwise covariate missingness patterns. However, in practice, missingness can occur anywhere within the dataset, e.g., for both outcomes and covariates under the regression setting.
@@
To address a more general missingness structure, we introduce the concept of missingness patterns and formalize the general missingness pattern problem by treating each missing pattern as a special case of coarsening (Heitjan and Rubin,, 1991), following Tsiatis, (2006). A missingness pattern is defined as an element in the sample space $\{0,1\}^{p}$ of the missing indicator vector $\bm{R}_{i}$. Excluding the case where all variables are missing, there are at most $2^{p}-1$ different missingness patterns appearing in the data. In the context of prediction-based inference, it is generally assumed that there exist fully observed, low-cost variables that serve as inputs for predicting the missing values (i.e., type (c) variables). Under this assumption, the number of distinct missingness patterns is reduced to at most $2^{q}$. Each of the missingness patterns can be indexed with $k\in\{1,...,K\}$; we define $k=\infty$ to indicate that the subject is fully observed. The $k$-th missingness pattern is denoted as $\bm{R}^{(k)}=(R^{(k)}_{1},\ldots,R^{(k)}_{p})\in\{0,1\}^{p}$. We call the $i$-th subject $\bm{Y}_{i}$ to be in the $k$-th stratum of missingness pattern if $\bm{R}_{i}=\bm{R}^{(k)}$. Following Tsiatis, (2006), for the $k$-th missingness pattern, we introduce a deterministic and many-to-one coarsening function $G_{k}:{\mathcal{Y}}\mapsto\prod_{\{j:R^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}$, such that $G_{k}(\bm{Y}_{i})=\bm{Y}_{i,\text{obs}}$ if $\bm{Y}_{i}$ is under the $k$-th missingness pattern. For fully observed data, $G_{\infty}(\bm{Y}_{i})$ is the identity mapping. Note that the coarsening notation is borrowed only to emphasize the constructive nature of our proposed method, which combines pattern-stratified estimates (see Equation (2)); the prediction-based inference for general coarsened data where $G_{k}$ is broader than entry subset selection is beyond the scope of this work. With the notation of the coarsening functions and missingness patterns, we can represent the observed dataset as ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$, where ${\mathcal{C}}_{i}\in\{1,...,K,\infty\}$ is the coarsening index for subject $i$.
@@
Missing data mechanisms play an important role in dictating how researchers should deal with the missingness in data analyses (Rubin,, 1976). Here we provide a brief overview for completeness and refer readers to Little and Rubin, (2019) for a comprehensive, book-length treatment of this topic. Based on how the conditional distribution of the coarsening index ${\mathcal{C}}_{i}$ given $\bm{Y}_{i}$ can be simplified, a few general categories with distinct analytic implications can be defined (Tsiatis,, 2006). In the simplest case with the most restrictive assumption, the probability distribution of the coarsening index ${\mathcal{C}}_{i}$ does not depend on any observed or unobserved variables, namely, $\pi({\mathcal{C}}_{i}=k|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i};{\bm{\omega}})$ where $\bm{\omega}$} is a vector of parameters of the conditional distribution. In this scenario, we consider the data to be missing completely at random (MCAR). A more common and realistic case compared to MCAR is that the missingness pattern depends on variables that are observed: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}};{\bm{\omega}})$, referred to as missing at random (MAR). If the missingness depends on the missing values themselves: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}},\bm{Y}_{i,\text{mis}};{\bm{\omega}}),$ it is referred to as missing not at random (MNAR). In practice, the MAR assumption is more commonly considered because it does not suffer from the identifiability problems that exists in the MNAR assumption, yet it is weaker than the MCAR assumption (Tsiatis,, 2006). In this paper, we will introduce and provide theoretical analyses of our methods under the MAR assumption.

### v2

@@ Abstract
Pre-trained machine learning (ML) predictions have been increasingly used to complement incomplete data to enable downstream scientific inquiries, but their naive integration risks biased inferences. Recently, multiple methods have been developed to provide valid inference with ML imputations regardless of prediction quality and to enhance efficiency relative to complete-case analyses. However, existing approaches are often limited to missing outcomes under a missing-completely-at-random (MCAR) assumption, failing to handle general missingness patterns (missing in both the outcome and exposures) under the more realistic missing-at-random (MAR) assumption. This paper develops a novel method that delivers a valid statistical inference framework for general Z-estimation problems using ML imputations under the MAR assumption and for general missingness patterns. The core technical idea is to stratify observations by distinct missingness patterns and construct an estimator by appropriately weighting and aggregating pattern-specific information through a masking-and-imputation procedure on the complete cases. We provide theoretical guarantees of asymptotic normality of the proposed estimator and efficiency dominance over weighted complete-case analyses. Practically, the method affords simple implementations by leveraging existing weighted complete-case analysis software. Extensive simulations are carried out to validate theoretical results. A real data example is provided to further illustrate the practical utility of the proposed method. The paper concludes with a brief discussion on practical implications, limitations, and potential future directions.
@@ 1 Introduction
With the rapid development of machine learning (ML) and artificial intelligence (AI) techniques, ML models, particularly those trained on large-scale datasets, have become increasingly capable of making highly accurate predictions. In statistical analysis, these predictions have the potential to be treated as imputations for missing data (in both dependent $Y$ and independent variables $\bm{X}$) due to nonresponse, expensive data collection costs, or other factors, and then the imputed datasets can be seamlessly applied in downstream statistical analyses. However, such an appealing impute-then-analyze procedure is a double-edged sword. On the one hand, the ease of obtaining ML predictions for imputation catalyzes wide adoption across various disciplines, including finance (bryzgalova2025missing), social science (egami2023using), clinical psychology (little2024missing), biomedicine (barrabes2024advances), and public health (perkins2018principled) studies, where missing data appear commonly during the data collection process. On the other hand, imputing missing values with ML predictions of arbitrary quality may produce datasets that have different conditional and marginal distributions from the underlying true distribution. Proceeding with analysis without careful consideration may result in biased estimation and invalid inference.
@@
To address this issue, recent research has explored prediction-based methods to incorporate ML predictions for single imputation while guaranteeing valid statistical inference on datasets with partially missing outcomes (see, for example, wang2020methods; angelopoulos2023prediction; egami2023using; salerno2025moment; salerno2025ipd). Specifically, prediction-powered inference (PPI, angelopoulos2023prediction) is a framework where inference is conducted using ML-imputed outcomes, with model parameter estimates calibrated, or “rectified”, using a complete-case subset. Since this approach does not impose assumptions on the imputation model itself, it allows for a broad selection of ML-based imputation methods, the quality of which impacts the efficiency but not the validity of inference (e.g., consistency, achieving nominal coverage rate). Several extensions of this framework have adapted the idea of PPI to important areas, including causal inference (demirel2024prediction), Bayesian inference (hofer2024bayesian; rovckova2025ai), and genome-wide association studies (GWAS, miao2024valid). Furthermore, recognizing that the PPI approach does not guarantee a smaller variance relative to complete-case analysis (CCA), a few authors have proposed weighting the rectifier with tuning parameters to achieve better efficiency than CCA (angelopoulos2023PPI++; miao2023assumption; gan2024prediction; gronsbell2024another; ji2025predictions; xu2025unified).
@@
Despite the tremendous progress and development in prediction-based inference, two key challenges remain unaddressed. First, the existing literature considers only the assumption of missing completely at random (MCAR), leaving an open question of the applicability of prediction-based inference under the more common and complicated assumption, that is, the assumption missing at random (MAR) (rubin1976inference; little2019statistical). Except for several specific cases such as linear regression analysis and odds ratio estimation under certain missing probability assumptions (Example 3.3 and 3.4, little2019statistical), failure to account for the MAR assumption can result in biased estimates. Second, the current literature focuses exclusively on partially missing outcomes or specific patterns in covariate missingness (e.g., missing multiple covariates kluger2025prediction), but does not account for general patterns of non-monotone missingness, where missing values can occur in any cell within the tabular (“rectangular”) dataset. For example, in regression analysis, missing values are common in both outcomes and covariates, and we may observe different patterns of missingness in different subjects, e.g., missing $Y$ and $X_{1}$ for one subject and missing $Y$ and $X_{2}$ for another.
@@
In this paper, we address these challenges by proposing a unified framework called Pattern-Stratified PPI (PS-PPI) [^1] By Corollary 3.2, the proposed procedure enjoys efficiency dominance over WCCA (under MAR, general missing patterns). This resembles the property of PPI++ being no less efficient than CCA (under MCAR, outcome missing only). We chose to term our method “PS-PPI” rather than “PS-PPI++” to avoid redundancy. to account for both general missingness patterns and missing-at-random mechanism under the Z-estimation framework. We make a few theoretical and methodological contributions with useful practical implications. First, we develop an inferential procedure that integrates the concept of missingness patterns that are used in semiparametric inference to deal with non-monotone missing data (tsiatis2006semiparametric). We will show that the proposed estimator belongs to the class of Augmented Inverse Probability Weighted (AIPW) estimators, building on a prior work under MCAR with only outcomes subject to missingness (gronsbell2024another). Second, we provide a theoretical guarantee that the resulting PS-PPI estimator is at least as efficient as the weighted complete-case analysis (WCCA) even when the imputation models are of low quality. Third, our proposed PS-PPI procedure is straightforward to implement. It works by breaking the analysis into several weighted complete-case (WCC) Z-estimation problems, each of which can be solved with existing statistical packages designed for WCC analysis.
@@
The remainder of the paper is organized as follows. In Section 2, we introduce necessary notation, the concepts of missing data mechanisms and missingness patterns, and the inferential target under the Z-estimation framework; we review the core idea used in the existing prediction-based inference literature that deals with missing outcome or missing multiple covariates, along with their limitations. In Section 3, we introduce the details of the proposed PS-PPI procedure and present theoretical results on its asymptotic properties; we provide examples to show that our proposed framework subsumes previous works as special cases. In Section 4, we perform data simulation to verify the validity and superior performance of our proposed PS-PPI relative to alternatives that have limited ability to extract and integrate information from incomplete datasets with multiple different missingness patterns. We provide a real data example in Section 5 to illustrate the practical utility of the PS-PPI method. Finally, we discuss the limitations of this work and potential future directions in Section 6.
@@ 2.1 Notation
Let $\mathcal{D}=\{\bm{Y}_{1},\ldots,\bm{Y}_{N}\}$ denote the full data for $N$ observations, subject to potential missingness, where $\bm{Y}_{i}=(Y_{i1},\ldots,Y_{ip})^{\top}\in{\mathcal{Y}}={\mathcal{Y}}_{1}\times\cdots\times{\mathcal{Y}}_{p}$ represents the $p$-dimensional data for subject $i$, ${\mathcal{Y}}_{j}\subseteq\mathbb{R}$ is the marginal sample space of the $j$-th component in $\bm{Y}_{i}$. Each $\bm{Y}_{i}$ is assumed to be independently and identically sampled from a population distribution $\mathbb{P}_{\bm{Y}}$. In practice, three types of variables may be included in the vector $\bm{Y}_{i}$: (a) a variable treated as the scalar outcome in a specific scientific inquiry; (b) q1q-1 variables treated as covariates (e.g., main exposure and potential confounders); (c) pqp-q auxiliary variables not included in the scientific analysis but potentially useful to predict variables in type (a) and (b). In a real data analysis, one or more variables of type (b) may not be subject to missingness, which can be used along with type (c) variables as inputs to ML prediction models for imputations. We partition $\bm{Y}_{i}$ into three types of variables for ease of presentation; this partition does not impact the validity of our theoretical results. For readers who require this distinction in their scientific inquiries (e.g., via a regression analysis), one may assume that $Y_{i1}$ and $Y_{i2},\ldots,Y_{iq}$} represent the outcome and the covariates, respectively.
We assume that the analyst may not observe the full data $\mathcal{D}$}. Let $\bm{R}_{i}=(R_{i1},\ldots,R_{ip})^{\top}\\ \in\{0,1\}^{p}$ be a binary vector of $p$ dimension, where $R_{ij}=1$ indicates that $Y_{ij}$ is observed and $R_{ij}=0$ otherwise. Consequently, we denote the observed components $\bm{Y}_{i,\text{obs}}=\{Y_{ij}|R_{ij}=1,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=1\}}{\mathcal{Y}}_{j}$, and similarly the missing components $\bm{Y}_{i,\text{mis}}=\{Y_{ij}|R_{ij}=0,j=1,\ldots,p\}\in\prod_{\{j:R_{ij}=0\}}{\mathcal{Y}}_{j}$, for subject $i=1,\ldots,N$. Let ${\mathcal{D}}_{\text{obs}}=\{\bm{Y}_{1,\text{obs}},...,\bm{Y}_{N,\text{obs}}\}$ denote the observed data.
@@ 2.1.1 Missingness patterns and mechanisms
In the existing literature on prediction-based inference, typically only the outcome variable is subject to potential missingness (e.g., angelopoulos2023prediction; angelopoulos2023PPI++; miao2023assumption; gan2023prediction; gronsbell2024another), with the exception of kluger2025prediction, which focused on the missing-multiple-covariate pattern. However, in practice, missingness can occur anywhere within the dataset, e.g., for both outcomes and covariates under the regression setting.
@@
To address a more general missingness structure, we use the concept of missingness patterns and formalize the general missingness pattern problem by treating each missingness pattern as a special case of coarsening (heitjan1991ignorability), following tsiatis2006semiparametric. A missingness pattern is defined as an element in the sample space $\{0,1\}^{p}$ of the missing indicator vector $\bm{R}_{i}$. Excluding the case where all variables are missing, there are at most $2^{p}-1$ different missingness patterns appearing in the data. In the context of prediction-based inference, it is generally assumed that there exist fully observed, low-cost variables that serve as inputs for predicting the missing values (i.e., type (c) variables). Under this assumption, the number of distinct missingness patterns is reduced to at most $2^{q}$. Each of the missingness patterns can be indexed with $k\in\{1,...,K\}$; we define $k=\infty$ to indicate that the subject is fully observed. The $k$-th missingness pattern is denoted as $\bm{R}^{(k)}=(R^{(k)}_{1},\ldots,R^{(k)}_{p})\in\{0,1\}^{p}$. We call the $i$-th subject $\bm{Y}_{i}$ to be in the $k$-th stratum of missingness pattern if $\bm{R}_{i}=\bm{R}^{(k)}$. Following tsiatis2006semiparametric, for the $k$-th missingness pattern, we introduce a deterministic and many-to-one coarsening function $G_{k}:{\mathcal{Y}}\mapsto\prod_{\{j:R^{(k)}_{j}=1\}}{\mathcal{Y}}_{j}$, such that $G_{k}(\bm{Y}_{i})=\bm{Y}_{i,\text{obs}}$ if $\bm{Y}_{i}$ is under the $k$-th missingness pattern. For fully observed data, $G_{\infty}(\bm{Y}_{i})$ is an identity mapping. Note that the coarsening notation is borrowed only to emphasize the constructive nature of our proposed method, which combines pattern-stratified estimates (see Equation (2)); the prediction-based inference for general coarsened data where $G_{k}$ is broader than entry subset selection is beyond the scope of this work. With the notation of the coarsening functions and missingness patterns, we can represent the observed dataset as ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$, where ${\mathcal{C}}_{i}\in\{1,...,K,\infty\}$ is the coarsening index for subject $i$.
@@
Missing data mechanisms play an important role in dictating how researchers should deal with the missingness in data analyses (rubin1976inference). Here we provide a brief overview for completeness and refer readers to little2019statistical for a comprehensive, book-length treatment of this topic. Based on how the conditional distribution of the coarsening index ${\mathcal{C}}_{i}$ given $\bm{Y}_{i}$ can be simplified, a few general categories with distinct analytic implications can be defined (tsiatis2006semiparametric). In the simplest case with the most restrictive assumption of missing completely at random (MCAR), the probability distribution of the coarsening index ${\mathcal{C}}_{i}$ does not depend on any observed or unobserved variables, namely, $\pi({\mathcal{C}}_{i}=k|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i};{\bm{\omega}})$ where $\bm{\omega}$} is a vector of parameters of the conditional distribution. A more common and realistic case compared to MCAR is that the probability of a missingness pattern only depends on observed variables: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}};{\bm{\omega}})$, referred to as missing at random (MAR). If the missingness depends on the missing values themselves: $\pi({\mathcal{C}}_{i}|\bm{Y}_{i};{\bm{\omega}})=\pi({\mathcal{C}}_{i}|\bm{Y}_{i,\text{obs}},\bm{Y}_{i,\text{mis}};{\bm{\omega}}),$ it is referred to as missing not at random (MNAR). In practice, the MAR assumption is more commonly considered because it does not suffer from the identifiability problems that exist in the MNAR assumption, yet it is weaker than the MCAR assumption (tsiatis2006semiparametric). In this paper, we will introduce and provide theoretical analyses of our methods under the MAR assumption.